OpenStudy (anonymous):

-6/7y -6 >= 42 Solve the inequality.

OpenStudy (anonymous):

Add both sides by 6, then divide both sides by -6/7 and change the inequality sign. Is it clear?

OpenStudy (anonymous):

yeah. thanks.

OpenStudy (anonymous):

No problem. Any question?

OpenStudy (anonymous):

nopee.

OpenStudy (anonymous):

wait, I do lol. How would you divide the 6/7 by 48?

OpenStudy (anonymous):

\[\Huge \frac{48}{\frac{6}{7}} = \frac{48}{\frac{6}{7}} \cdot \frac{7}{7} = \frac{48 \cdot 7}{\frac{6}{7} \cdot 7} = \frac{48 \cdot 7}{6}\] Is this clear?

OpenStudy (anonymous):

yes.